OpenStudy (anonymous):

how to factor x^2 - 7x + 12

OpenStudy (anonymous):

Lets do this question by grouping. so let's separate the 7x in the middle into -3x-4x x^2 - 7x + 12 x^2 - 3x - 4x + 12 can you do the next step of grouping?

OpenStudy (anonymous):

12= 4 x 3. -4-3 = -7. -4 X -3 = +12 (x-4)(x-3) hence factrs r 4 and 3.

OpenStudy (anonymous):

I guess dis s a way f teachin! sry

OpenStudy (anonymous):

Hope u learnt Mr. plohrr!

OpenStudy (anonymous):

i get it thanks !